Question 1.)

a.) Evaluate a 5-year project costing $32,000 and returning $5,000 annually using the payback period technique and a 4-year cutoff.

b.) Evaluate the project above using the NPV method and a 13% required rate.

c.) Evaluate the project in part "a" using the profitability index method and a 10% required return.

Please show formulas used (not a spreadsheet) and how you worked out the problem, thank you!
Answer #1

a: The cumulative cash flows are

Year 1 = -32000+5000 = -27000

Year 2 = -27000+5000 = -22000

Year 3 = -22000+5000 = -17000

Year 4 = -17000+5000=-12000

Year 5 = -12000+5000 = -7000

The project cannot be accepted since investment is not recovered till cut off period of year 4

b: NPV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= -32000+5000/1.13^1+5000/1.13^2+5000/1.13^3+5000/1.13^4+5000/1.13^5=

-14413.84

Reject the project since NPV is negative

c: PI = PV of inflows/Initial cost = 5000/1.1^1+5000/1.1^2+5000/1.1^3+5000/1.1^4+5000/1.1^5/32000

=0.59

Reject the project since Profitability index is less than 1